1. Solve the differential equation. Use the fact that the given equation is homogeneous.
dy / dx = (x^2 + 8y^2) / 3xy

2. A certain college graduate borrows 7486 dollars to buy a car. The lender charges interest at an annual rate of 19% . Assuming that interest is compounded continuously and that the borrower makes payments continuously at a constant annual rate k dollars per year, determine the payment rate that is required to pay off the loan in 2 years. Also determine how much interest is paid during the 2-year period.

Round your answers to two decimal places.

A. Payment rate ______ dollars per year. (absolute tolerance +/- 0.01)
B. Interest Paid _____ dollars (absolute tolerance is +/- 0.05)
1
Expert's answer
2020-06-15T19:36:32-0400

1) "\frac {dy}{dx}" = "\frac {x²+8y²}{3xy}" = "\frac {1+8(y/x)²}{3y/x}"

Since "\frac {dy}{dx}" is a function of y/x , it is homogeneous form.

To solve this

Put y = vx

Differentiating "\frac {dy}{dx}" = v + x"\frac {dv}{dx}"

So v + "x\frac {dv}{dx}" = "\frac {1+8(v)²}{3v}"

=> "x\frac {dv}{dx}" = "\frac {1+8(v)²}{3v}" - v = "\frac {1+5v²}{3v}"

=> "\frac {vdv}{5v²+1} = \frac {dx}{3x}"

=> "\frac {vdv}{v²+\frac {1}{5}} = \frac {5}{3}\frac {dx}{x}"

Integrating,

"\int\frac {vdv}{v²+\frac {1}{5}} =\int \frac {5}{3}\frac {dx}{x}"

Let v²+"\frac {1}{5}" = z

So 2 v dv = dz

"\frac {1}{2}\int\frac {dz}{z} =\ \frac {5}{3}\int\frac {dx}{x}"

"3\int\frac {dz}{z} =10\int\frac {dx}{x}"

=> 3 ln |z |= 10 ln | x| + lnA

=> ln|z|³= ln x10+lnA

=> ln |z|³ = ln (Ax10)

=> |z|³ = Ax10

=> |v²+ "\frac {1}{5}" |³ = Ax10

=> | 5v²+1|³ = 125Ax10

=> | 5y²+ x²|³ = 125Ax6x10

=> (5y²+x²)³ = C³x16 Considering 125A = C³

=> (x²+5y²) = C"x^{\frac {16}{3}}"

This is the general solution of given differential equation


2) Let x be the amount borrowed by the student and t represents time

So "\frac {dx}{dt}" = "\frac {19x}{100} - k" = 0.19x - k

=> "\frac {dx}{0.19x-k} = dt"

Integrating,

"\int\frac {dx}{0.19x-k} = \int dt"

"\frac {1}{0.19}" ln |0.19x-k| = t + C

As the loan is repayable , 0.19x < k

So ln (k-0.19x) = 0.19t + 0.19C

When t = 0, x = 7486

So 0.19C = ln(k - 0.19*7486)

=> ln (k-0.19x) = 0.19t + ln(k - 0.19*7486)

As the borrower wants to pay off the loan in 2 years, when t = 2 , x = 0

So ln k = 0.38 + ln(k - 0.19*7486)

=> ln "\frac { k} { k - 0.19*7486}" = 0.38

=> k = (k - 0.19*7486)e0.38

=> k(e0.38-1) = 0.19*7486*e0.38

=> k = "\frac {0.19*7486*e^{0.38}}{e^{0.38}-1}"

=> k = 4499.10

As loan is repaying at a constant rate of 4499.10 dollars per year, inv2-year period borrower pays 2*4499.10 dollars = 8998.20 dollars.

So interest in 2-year period

= Amount paid in 2 years subtracted by amount of loan taken

= 8998.20 - 7486 dollars

= 1512.20 dollars

Answer

A. Payment rate 4499.10 dollars per

year

B. Interest paid 1512.20 dollars